25j of energy is stored in a capacitor ot capacitance 2NF . Determine the voltage across the capacitor terminal


1
Expert's answer
2021-08-16T00:53:02-0400

"E=1\/2CV^2"

"2E=CV^2"

"V^2=2E\/C"

"V=\sqrt{2E/C}" ="\sqrt{\frac{(2\times 25)}{2\times 10^{-9}}}=158113.883" "V"